# Quantum mixing on large Schreier graphs

### v1

@@ 6.3 Lifts of a finite base graph
Moreover, setting $\zeta^{z}_{i}=-a_{i}\hat{\gamma}_{i}^{z}$, since $p_{g_{i}}=a_{i}E_{u_{i}v_{i}}$, the entry $(u,v)$ of $R^{z}(e,g)$ in (6.17) is equal to
$R^{z}(e,e)_{uu_{i_{1}}}\zeta_{i_{1}}^{z}\ldots\zeta_{i_{\ell}}^{z}$ (6.19)
if $v_{i_{j}}=u_{i_{j+1}}$ for any $1\leq j\leq\ell-1$ and $v_{i_{\ell}}=v$, and is equal to 0 otherwise. Also, for all but a finite set $D$ if $\Re(z)\notin D$, for all $u\in\left[r\right]$, the matrix entries $R^{z}(e,e)_{uu}$ have non-vanishing modulus in the limit $\eta\rightarrow 0$ (this is mostly due to [7, Proposition 4.2]). Therefore, the Ward identity implies that
@@
for some $C>0$ independent of $\eta$, where $\mathcal{I}_{\ell}$ is the set of non-backtracking paths of length $\ell$ in the base graph $H$:
$\mathcal{I}_{\ell}=\{(i_{1},\ldots,i_{\ell})\in[2d]^{\ell}\mid i_{k}\neq i_{i_{k+1}}^{*},v_{i_{k}}=u_{i_{k+1}}\text{\ for any\ }1\leq k\leq\ell-1\}.$
Since $R^{z}(e,e)$ has all its entries uniformly bounded above in modulus as $\eta\rightarrow 0$, proving (5.5) reduces to showing
@@
Such a bound essentially follows from [53, Sec. 5.2] in case the minimal degree is $\geq 3$. We provide below an alternative argument which does not need this restriction.
We call a loop any sequence of the form $(i_{1},\ldots,i_{\ell})\in\mathcal{I}_{\ell}$ for some $\ell\in\mathbb{N}$, such that $v_{i_{\ell}}=u_{i_{1}}$, $i_{1}\neq i_{\ell}^{*}$, and all the $i_{j}$ are distinct (note that the visited vertices $(u_{i_{1}},\ldots,u_{i_{\ell}})$ are not necessarily distinct). Two loops are distinct if they are not equal up to cyclic ordering. Our assumption on $H$ implies that for any loop $(i_{1},\ldots,i_{\ell})$, there exists a distinct loop $(j_{1},\ldots,j_{k})$ sharing a common vertex. We may assume without loss of generality that $u_{i_{1}}=u_{j_{1}}$. Moreover, up to replacing $(j_{1},\ldots,j_{k})$ by its inverse $(j_{k}^{*},\ldots,j_{1}^{*})$, we may assume that the two loops can be composed as non-backtracking paths (note one loop could be a subloop of the other, for example if $H$ is a dumbbell graph). Then, repeating successively these two loops in any order, (6.20) implies that
@@
$\qquad\qquad\leq C\eta\sup_{\ell\in\mathbb{N}}\left(\ell^{C_{1}^{\prime}}{c_{0}^{\prime}}^{2}(1-\delta^{\prime})^{2\ell}\right).$
This concludes the proof of (6.21), and thus of (5.5). ∎
@@ 7 Necessity of the assumptions
We take the same example as in [61, Section 4]. Consider any family of $d$-regular graphs $(F_{N})$ for $d$ even and $d\geq 8$ such that $|F_{N}|=N$ and $(F_{N})$ converges in the sense of Benjamini-Schramm to the $d$-regular tree. We construct a family of graphs $(G_{N})$ as follows (see [61, Figure 3]). First, we delete an arbitrary edge of $F_{N}$, and call this new graph $F_{N}^{\prime}$. We create $d/2$ copies of $F_{N}^{\prime}$, add a vertex $v_{N}$, and add an edge from $v_{N}$ to each of the $d$ vertices of degree d1d-1, two for each copy of $F_{N}^{\prime}$. We call this graph $G_{N}$, and enumerate the copies of $F_{N}^{\prime}$ in $G_{N}$ as $F_{N,1}^{\prime},\ldots,F_{N,d/2}^{\prime}$. Since $G_{N}$ is a $d$-regular graph of even degree, it is a Schreier graph, and its adjacency matrix is of the form $\rho_{N}(\mathbf{1}_{S})$ for some representation $\rho_{N}$ and for the canonical set $S$ of generators of the free group $\mathbf{F}_{d}$ (see Section 1.1). Moreover, $\rho_{N}$ converges weakly in distribution to the regular representation of $\mathbf{F}_{d}$, since $(F_{N})$ (and thus $(G_{N})$) converges in the Benjamini-Schramm sense to the $d$-regular tree.
@@
$a_{N}(x)=\begin{cases}1&\text{if }x\in V_{F^{\prime}_{N,1}}\cup V_{F^{\prime}_{N,2}}\\ -1&\text{if }x\in V_{F^{\prime}_{N,3}}\cup V_{F^{\prime}_{N,4}}\\ 0&\text{otherwise.}\end{cases}$
@@ 7.2.1 The $\mathbb{Z}^{d}$ case: no quantum mixing
  • the spectral measure of the adjacency matrix of the Cayley graph $\mathbb{Z}^{d}$ is obtained as the convolution of $d$ arcsine distributions. Hence it is absolutely continuous, and with density bounded above and below on any compact subset of its support (2d,2d)(-2d,2d), in particular (1.5) holds on any compact interval $I_{0}\subset(-2d,2d)$;
@@
Hence, if $I_{\eta}$ is an interval of size $\eta$ included in the interior of (2d,2d)(-2d,2d) and $\langle a\rangle=0$ (meaning that $\widehat{a}(0)=0$), we have
$\frac{1}{|\Lambda_{I_{\eta}}|}\sum_{\lambda_{k},\lambda_{k^{\prime}}\in I_{\eta}}|\langle\varphi_{k},a\varphi_{k^{\prime}}\rangle|^{2}=\frac{1}{|\Lambda_{I_{\eta}}|}\sum_{\lambda_{k}\in I_{\eta}}\sum_{\lambda_{k^{\prime}}\in I_{\eta}}\frac{|\widehat{a}(k^{\prime}-k)|^{2}}{N}.$ (7.1)
@@
We claim that if $\eta>0$, for $N$ large enough, at least $\frac{1}{2}|\Lambda_{I_{\eta}}|$ eigenvalues $\lambda_{k}\in I_{\eta}$ satisfy $\lambda_{k+u}\in I_{\eta}$. Let us prove this claim. Since $I_{\eta}$ is included in the interior of (2d,2d)(-2d,2d), the spectral density in this interval of the Cayley graph $\mathbb{Z}^{d}$ with standard generators is bounded and strictly positive (as a convolution of arcsine densities). Using that the spectral measure of the box $(\mathbb{Z}/M\mathbb{Z})^{d}$ converges weakly as $M\rightarrow+\infty$ to the spectral measure of the Cayley graph of $\mathbb{Z}^{d}$, and using also the fact that $|\lambda_{k+u}-\lambda_{k}|\leq\frac{2\pi}{M}\rightarrow 0$ as $M\rightarrow+\infty$, we get the claim.
We now conclude the proof of Proposition 7.2 using the above claim. We choose $a(n)=e^{\frac{2i\pi n_{1}}{M}}$, i.e., $\widehat{a}(\ell)=M^{d/2}\delta_{\ell=u}$. Then the sum in (7.1) is greater than $\frac{1}{2}$, which shows that quantum mixing (1.11) does not hold.
@@
In this section, we provide a family of graphs in which all assumptions of Theorem 5.3(iii) hold, except the $4$-moment condition (5.5) on the resolvent, and quantum ergodicity in the sense of (1.8) fails. This shows the necessity of a resolvent assumption like (5.5) for quantum ergodicity to hold. This example was first considered in [62].
@@
$\frac{1}{|\Gamma_{N}|}\sum_{u\in\Gamma_{N}}|\langle\psi_{u},a_{N}\psi_{u}\rangle|^{2}\geq\frac{1}{320}.$
@@
In this section, we provide an example showing the necessity of the assumption $a_{N}\in\mathbb{C}^{r}\otimes 1^{\perp}$ in Theorem 5.3(iii); in particular assuming $\langle a_{N}\rangle=0$ would not be sufficient to get Theorem 5.3(iii). The family of graph considered below has also been used by McKenzie in [61], who showed that in this family of graphs, quantum ergodicity fails for some observables. Our Theorem 5.3(iii) shows that quantum ergodicity holds nevertheless in the class of observables $\mathbb{C}^{r}\otimes 1^{\perp}$.
In the sequel, $d$ is even. Let $g_{1},\ldots,g_{d/2},g_{1}^{-1},g_{d/2}^{-1}$ be the set of generators of the free group on $d$ generators $\mathbf{F}_{d}$, and let
@@
where $\rho_{N}$ is a random representation of $\mathbf{F}_{d}$ in the symmetric group $S_{N}$. According to [15, Theorem 3], $\rho_{N}$ converges strongly (in probability) to the regular representation $\lambda$.
We consider the Cartesian product $C_{4}\Box G_{N}$ where $G_{N}$ is the $d$-regular graph with adjacency matrix $A_{N}$, and $C_{4}$ is the cycle graph with $4$ elements. Its adjacency matrix $A_{C_{4}}\otimes I_{G_{N}}+I_{C_{4}}\otimes A_{N}$ is of the form (5.2) with $p_{e}=A_{C_{4}}$ the adjacency matrix of the $4$-cycle, and $p_{g}=I_{C_{4}}$ for $g\in S$ where $\Gamma$ is the free group generated by $S$, with $d/2$ elements and their inverses.
The strong convergence condition in Theorem 5.3(iii) holds, by our choice of $\rho_{N}$. The rapid decay condition is also satisfied in the $d$-regular tree (see [26]). Since the spectrum of the $d$-regular tree $\mathbb{T}_{d}$ is absolutely continuous, the spectral measure of $C_{4}\Box\mathbb{T}_{d}$, which is the convolution of the spectral measures of $C_{4}$ and $\mathbb{T}_{d}$, is absolutely continuous too. Finally, the condition (5.5) is satisfied, as a consequence of [61, Theorem 1.3]. Theorem 5.3(iii) thus applies and shows that quantum ergodicity holds in this family of graphs for the class of observables $\mathbb{C}^{r}\otimes 1^{\perp}$.
However, the conclusion of Theorem 5.3(iii) fails in the family of graphs $C_{4}\Box G_{n}$ if one considers observables which are not in $\mathbb{C}^{r}\otimes 1^{\perp}$: as shown in [61], there exists for any $n$ an orthonormal eigenbasis $(\psi_{j})$ of the adjacency matrix of $C_{4}\Box G_{n}$ and an observable $a_{n}$ with $\langle a_{n}\rangle=0$ and $\|a_{n}\|_{\infty}=1$ such that
@@
Let us briefly describe the eigenbasis and the observables. The eigenvectors $\psi_{j}$ are taken in a “separation of variables” form, namely $\psi_{j}(i,x)=\phi_{1}(i)\phi_{2}(x)$ for $(i,x)$ a vertex in the graph $C_{4}\Box G_{n}$, with $\phi_{1},\phi_{2}$ eigenvectors of the adjacency matrices of $C_{4}$ and $G_{n}$ respectively. The eigenvectors of $C_{4}$ that we take for $\phi_{1}$ are
@@
We define the observable $a_{n}$ by its values at $(i,x)\in C_{4}\times G_{n}$: $a_{n}(i,x)=1$ for $i\in\{1,3\}$ and $a_{n}(i,x)=-1$ for $i\in\{2,4\}$. It is not difficult to check that (7.2) holds, see [61, Section 2].
@@
Theorem 5.3(iii) implies nevertheless that (5.10) holds. The above array of eigenvectors of $C_{4}$ shows that the quantity $\|\varphi_{\alpha}(i,\cdot)\|^{2}$ appearing in (5.10) depends on the eigenvector $\varphi_{\alpha}$. In fact, taking the $C_{4}$ eigenvectors $(\frac{1}{2},\frac{e^{ik}}{2},\frac{e^{2ik}}{2},\frac{e^{3ik}}{2})$, $k=0,1,2,3$ would not yield a counterexample.

### v2

@@ 6.3 Lifts of a finite base graph
The matrix $R^{z}(e,e)\in M_{r}(\mathbb{C})$ is diagonal due to (6.19) and the fact that for any sequence $(j_{1},\ldots,j_{k})$ such that $g_{j_{k}}\ldots g_{j_{1}}=e$, we have $u_{j_{1}}=v_{j_{k}}$. Let $\zeta^{z}_{i}=-a_{i}\hat{\gamma}_{i}^{z}$. Since $g=g_{i_{\ell}}\ldots g_{i_{1}}$ and $p_{g_{i}}=a_{i}E_{v_{i}u_{i}}$, the entry $(v,u)$ of $R^{z}(g,e)$ in (6.18) is equal to
$R^{z}(g,e)_{vu}=\zeta_{i_{\ell}}^{z}\ldots\zeta_{i_{1}}^{z}R^{z}(e,e)_{u_{i_{1}}u}$ (6.21)
if $u_{i_{1}}=u$, $v_{i_{\ell}}=v$, and $v_{i_{j}}=u_{i_{j+1}}$ for any $1\leq j\leq\ell-1$. It is equal to 0 otherwise. Hence, the Ward identity (2.16) becomes in the present context
$\Im R^{z}(e,e)_{uu}=\sum_{g,v}\eta|R^{z}(g,e)_{vu}|^{2}=\eta\sum_{\ell}\sum_{\begin{subarray}{c}(i_{1},\ldots,i_{\ell})\in\mathcal{I}_{\ell}\\ u_{i_{1}}=u\end{subarray}}|\zeta_{i_{1}}^{z}\ldots\zeta_{i_{\ell}}^{z}|^{2}|R^{z}(e,e)_{u_{i_{1}}u}|^{2}.$ (6.22)
Here,
$\mathcal{I}_{\ell}=\{(i_{1},\ldots,i_{\ell})\in[2d]^{\ell}\mid i_{k}\neq i_{k+1}^{*},v_{i_{k}}=u_{i_{k+1}}\text{\ for any\ }1\leq k\leq\ell-1\},$
that is, each $i\in\mathcal{I}_{\ell}$ indexes a sequence of $\ell$ adjacent directed edges $(u_{1},v_{1})$, $(u_{2},v_{2})$, etc. in $H$ without backtracking. For all but a finite set $D$ if $\Re(z)\notin D$, for all $u\in\left[r\right]$, the matrix entries $R^{z}(e,e)_{uu}$ have non-vanishing modulus in the limit $\eta\rightarrow 0$ (this is mostly due to [7, Proposition 4.2]). Therefore, (6.22) implies that
@@
for some $C>0$ independent of $\eta$.
Step 2. In view of (6.21) and since $R^{z}(e,e)$ has all its entries uniformly bounded above in modulus as $\eta\rightarrow 0$, proving (5.5) reduces to showing
@@
Such a bound essentially follows from [53, Section 5.2] in case the minimal degree is $\geq 3$. We provide below an alternative argument which does not need this restriction.
We call a loop any sequence of the form $(i_{1},\ldots,i_{\ell})\in\mathcal{I}_{\ell}$ for some $\ell\in\mathbb{N}$, such that $v_{i_{\ell}}=u_{i_{1}}$, $i_{1}\neq i_{\ell}^{*}$, and all the $i_{j}$ are distinct (note that the visited vertices $(u_{i_{1}},\ldots,u_{i_{\ell}})$ are not necessarily distinct). Two loops are distinct if they are not equal up to cyclic ordering. Our assumption on $H$ implies that for any loop $(i_{1},\ldots,i_{\ell})$, there exists a distinct loop $(j_{1},\ldots,j_{k})$ sharing a common vertex. We may assume without loss of generality that $u_{i_{1}}=u_{j_{1}}$. Moreover, up to replacing $(j_{1},\ldots,j_{k})$ by its inverse $(j_{k}^{*},\ldots,j_{1}^{*})$, we may assume that the two loops can be composed as non-backtracking paths (note that one loop could be a subloop of the other, for example if $H$ is a dumbbell graph). Then, repeating successively these two loops in any order, (6.23) implies that
@@
$\qquad\qquad\leq C\eta\sup_{\ell\in\mathbb{N}}\left(\ell^{C_{1}^{\prime}}{c_{0}^{\prime}}^{2}(1-\delta^{\prime})^{2\ell}\right)\longrightarrow 0$
as $\eta\to 0$. This concludes the proof of (6.24), and thus of (5.5). ∎
@@ 7 Necessity of the assumptions
We take the same example as in [61, Section 4]. Consider any family of $d$-regular graphs $(H_{N})$ for $d$ even and $d\geq 8$ such that $|H_{N}|=N$ and $(H_{N})$ converges in the sense of Benjamini-Schramm to the $d$-regular tree. We construct a family of graphs $(G_{N})$ as follows (see [61, Figure 3]). First, we delete an arbitrary edge of $H_{N}$, and call this new graph $H_{N}^{\prime}$. We create $d/2$ copies of $H_{N}^{\prime}$, add a vertex $v_{N}$, and add an edge from $v_{N}$ to each of the $d$ vertices of degree d1d-1, two for each copy of $H_{N}^{\prime}$. We call this graph $G_{N}$, and enumerate the copies of $H_{N}^{\prime}$ in $G_{N}$ as $H_{N,1}^{\prime},\ldots,H_{N,d/2}^{\prime}$. Since $G_{N}$ is a $d$-regular graph of even degree, it is a Schreier graph, and its adjacency matrix is of the form $\rho_{N}(\mathbf{1}_{S})$ for some representation $\rho_{N}$ and for the canonical set $S$ of generators of the free group $F_{d}$ (see Section 1.1). Moreover, $\rho_{N}$ converges weakly in distribution to the regular representation of $F_{d}$, since $(H_{N})$ (and thus $(G_{N})$) converges in the Benjamini-Schramm sense to the $d$-regular tree.
@@
$a_{N}(x)=\begin{cases}1&\text{if }x\in V_{H^{\prime}_{N,1}}\cup V_{H^{\prime}_{N,2}}\\ -1&\text{if }x\in V_{H^{\prime}_{N,3}}\cup V_{H^{\prime}_{N,4}}\\ 0&\text{otherwise.}\end{cases}$
@@ 7.2.1 The $\mathbb{Z}^{d}$ case: no quantum mixing
  • the spectral measure of the adjacency matrix of the Cayley graph $\mathbb{Z}^{d}$ is obtained as the convolution of $d$ arcsine distributions. Its support is [2d,2d][-2d,2d] and is absolutely continuous. For $d=2$, the density is bounded above and below except at $\{-4,0,4\}$, in particular (1.5) holds on any compact interval $I_{0}\subset(-4,4)\setminus\{0\}$. For $d\geq 3$ the density is bounded over (2d,2d)(-2d,2d) and (1.5) holds on any compact interval $I_{0}\subset(-2d,2d)$.
@@
Hence, fixing $\eta>0$, if $I^{\eta}$ is an interval of size $\eta$ included in the interior of (2d,2d)(-2d,2d) and not containing 0 if $d=2$, and $\langle a\rangle=0$ (meaning that $\widehat{a}(0)=0$), we have
$\frac{1}{|\Lambda_{I^{\eta}}|}\sum_{k,k^{\prime}\in\Lambda_{I^{\eta}}}|\langle\varphi_{k},a\varphi_{k^{\prime}}\rangle|^{2}=\frac{1}{|\Lambda_{I^{\eta}}|}\sum_{k\in\Lambda_{I^{\eta}}}\sum_{k^{\prime}\in\Lambda_{I^{\eta}}}\frac{|\widehat{a}(k^{\prime}-k)|^{2}}{N}.$ (7.1)
@@
We claim that for $N$ large enough, at least $\frac{1}{2}|\Lambda_{I^{\eta}}|$ eigenvalues $\lambda_{k}\in I^{\eta}$ satisfy $\lambda_{k+u}\in I^{\eta}$. In fact, if $I^{\eta}=[b,c]$, choose $I_{0}^{\delta}:=[b+\delta,c-\delta]\subset I^{\eta}$ such that $\mu_{p}(I_{0}^{\delta})\geq\frac{3}{4}\mu_{p}(I^{\eta})$, where $\mu_{p}$ is the spectral density of $\mathbb{Z}^{d}$ with standard generators. This choice is possible for $\delta>0$ small enough since $\mu_{p}$ has a bounded strictly positive density on $I^{\eta}$ by our assumptions. Since the spectral measure of the box $(\mathbb{Z}/M\mathbb{Z})^{d}$ converges weakly as $M\rightarrow+\infty$ to $\mu_{p}$, we have $|\Lambda_{I^{\delta}_{0}}|\geq\frac{1}{2}|\Lambda_{I^{\eta}}|$ for $M$ large enough. Using finally the fact that $|\lambda_{k+u}-\lambda_{k}|\leq\frac{2\pi}{M}\leq\delta$ for $M$ large enough, we get that all eigenvalues $\lambda_{k}$ in $I^{\delta}_{0}$ satisfy $\lambda_{k+u}\in I^{\eta}$, which concludes the proof of the claim.
We now conclude the proof of Proposition 7.2. We choose $a(n)=e^{\frac{2i\pi n_{1}}{M}}$, i.e., $\widehat{a}(\ell)=M^{d/2}\delta_{\ell=u}$. By the previous paragraph, at least half the $k\in\Lambda_{I^{\eta}}$ satisfy $\sum_{k^{\prime}\in\Lambda_{I^{\eta}}}\frac{|\widehat{a}(k^{\prime}-k)|^{2}}{N}\geq\frac{|\widehat{a}(k+u-k)|^{2}}{N}=1$. Hence, the sum in (7.1) is greater than $\frac{1}{2}$, which shows that quantum mixing (1.11) does not hold.
@@
In this section, we provide a family of graphs in which all assumptions of Theorem 5.3(iii) hold, except the $4$-th moment condition (5.5) on the resolvent, and quantum ergodicity in the sense of (1.8) fails. This shows the necessity of a resolvent assumption like (5.5) for quantum ergodicity to hold. This example was first considered in [62].
@@
$\frac{1}{|\Gamma_{N}|}\sum_{j\in\Gamma_{N}}|\langle\psi_{j},a_{N}\psi_{j}\rangle|^{2}\geq\frac{1}{320}.$
@@
In this section, we provide an example showing the necessity of the assumption $a_{N}\in\mathbb{C}^{r}\otimes 1^{\perp}$ in Theorem 5.3(iii); in particular assuming $\langle a_{N}\rangle=0$ would not be sufficient to get Theorem 5.3(iii). The family of graphs considered below has also been used by McKenzie in [61], who showed that in this family, quantum ergodicity fails for some observables. Our Theorem 5.3(iii) shows that quantum ergodicity holds nevertheless in the class of observables $\mathbb{C}^{r}\otimes 1^{\perp}$.
In the sequel, $d$ is even. Let $g_{1},\ldots,g_{d/2},g_{1}^{-1},g_{d/2}^{-1}$ be the set of generators of the free group on $d$ generators $F_{d}$, and let
@@
where $\rho_{N}$ is a random representation of $F_{d}$ in the symmetric group $S_{N}$. According to [15, Theorem 3], $\rho_{N}$ converges strongly (in probability) to the regular representation $\lambda$.
We consider the Cartesian product $C_{4}\mathop{\Box}G_{N}$ where $G_{N}$ is the $d$-regular graph with adjacency matrix $A_{N}$, and $C_{4}$ is the cycle graph with $4$ elements. Its adjacency matrix $A_{C_{4}}\otimes I_{G_{N}}+I_{C_{4}}\otimes A_{N}$ is of the form (5.2) with $p_{e}=A_{C_{4}}$ the adjacency matrix of the $4$-cycle, and $p_{g}=I_{C_{4}}$ for $g\in S$ where $\Gamma$ is the free group generated by $S$, with $d/2$ elements and their inverses.
The strong convergence condition in Theorem 5.3(iii) holds, by our choice of $\rho_{N}$. The rapid decay condition is also satisfied in the $d$-regular tree (see [26]). Since the spectrum of the $d$-regular tree $\mathbb{T}_{d}$ is absolutely continuous, the spectral measure of $C_{4}\mathop{\Box}\mathbb{T}_{d}$, which is the convolution of the spectral measures of $C_{4}$ and $\mathbb{T}_{d}$, is absolutely continuous too. Finally, the condition (5.5) is satisfied, as a consequence of [61, Theorem 1.3]. Theorem 5.3(iii) thus applies and shows that quantum ergodicity holds in this family of graphs for the class of observables $\mathbb{C}^{r}\otimes 1^{\perp}$.
However, the conclusion of Theorem 5.3(iii) fails in the family of graphs $C_{4}\mathop{\Box}G_{n}$ if one considers observables which are not in $\mathbb{C}^{r}\otimes 1^{\perp}$: as shown in [61], there exists for any $n$ an orthonormal eigenbasis $(\psi_{j})$ of the adjacency matrix of $C_{4}\mathop{\square}G_{n}$ and an observable $a_{n}$ with $\langle a_{n}\rangle=0$ and $\|a_{n}\|_{\infty}=1$ such that
@@
Let us briefly describe the eigenbasis and the observables. The eigenvectors $\psi_{j}$ are taken in a “separation of variables” form, namely $\psi_{j}(i,x)=\phi_{1}(i)\phi_{2}(x)$ for $(i,x)$ a vertex in the graph $C_{4}\mathop{\Box}G_{n}$, with $\phi_{1},\phi_{2}$ normalized eigenvectors of the adjacency matrices of $C_{4}$ and $G_{n}$ respectively. The eigenvectors $\phi_{1}$ of $C_{4}$ are chosen as follows:
@@
Define the observable $a_{n}$ by its values at $(i,x)\in C_{4}\times G_{n}$: $a_{n}(i,x)=1$ for $i\in\{1,3\}$ and $a_{n}(i,x)=-1$ for $i\in\{2,4\}$. It is not difficult to check that (7.2) holds, see [61, Section 2].
@@
Theorem 5.3(iii) implies nevertheless that (5.10) holds. The above array of eigenvectors of $C_{4}$ shows that the quantity $\|\varphi_{\alpha}(i,\cdot)\|^{2}$ appearing in (5.10) depends on the eigenvector $\varphi_{\alpha}$. In fact, taking the $C_{4}$ eigenvectors $(\frac{1}{2},\frac{e^{ik}}{2},\frac{e^{2ik}}{2},\frac{e^{3ik}}{2})$, $k=0,1,2,3$ would not yield a violation of ergodicity.